2 FeCl2 + Cl2 → 2 FeCl3

If 2.4154 moles of FeCl2 are reacted, how many moles of Cl2 will also be reacted?


1
Expert's answer
2021-02-02T01:23:26-0500

Solution:

The balanced chemical equation:

2FeCl2 + Cl2 → 2FeCl3

According to the equation above: n(FeCl2)/2 = n(Cl2)

Hence,

n(Cl2) = n(FeCl2)/2 = 2.4154 mol / 2 = 1.2077 mol

n(Cl2) = 1.2077 mol


Answer: 1.2077 moles of Cl2 will also be reacted.